OpenStudy (anonymous):

The energy density of an electromagnetic wave is ε0*E^2. To calculate the energy flux, at least in the derivation's I've seen, people just multiply by the speed of the wave, i.e., c. But doesn't this assume that the energy density is constant at all points, but E changes periodically. Why isn't it then the integral of the energy density in that volume, so it would give something close to a half of the usual answer i see.

OpenStudy (anonymous):

it's S=ε0*Ε^2*c/2

OpenStudy (anonymous):

that is for both the magnetic and the electric field...ε0*Ε/2 is if you take the contribution of only the electric field

OpenStudy (anonymous):

Just read in here: http://hep.ph.liv.ac.uk/~hutchcroft/...CN6EMWaves.pdf that I was right. ε0*E^2 is only the instantanous flux, but normally you would average it and get a factor of a half.

OpenStudy (anonymous):

link isnt working

OpenStudy (anonymous):

your energy is for an electric wave not an em wave

OpenStudy (anonymous):

electric field*

OpenStudy (anonymous):

http://hep.ph.liv.ac.uk/~hutchcroft/Phys258/CN6EMWaves.pdf Sorry here it is

OpenStudy (anonymous):

When the total instantaneous energy density is averaged over one or more cycles of an em wave you obtain a factor of 1/2 so yes you are right

OpenStudy (anonymous):

instantaneous energy for electric only=1/2ε0*Ε^2 total instantaneous of an em wave=electric+magnetic=ε0E^2 average energy density(total average density)=1/2ε0*Ε^2

OpenStudy (anonymous):

sorry forgot to write average density of the EM wave

OpenStudy (anonymous):

mutliplied by time? isn't the power flux equal to the energy density multiplied by the velocity?

OpenStudy (anonymous):

yes average energy density multiplied by velocity